a)     A simple closed economy, with an mpc equal to 0.75. The government has passed a balanced budget amendment. The economy goes into a recession, so the government increases government spending by 40 million to try to expand the economy

                   i.           Find the change in output from the increase in government spending.

                 ii.           The balanced budget amendment requires the government to also raise taxes by 40 million. Calculate the change in output from the tax hike.


1
Expert's answer
2021-07-05T16:25:38-0400

(i) The change in output from increase in government spending can be calculated with the help of income spending multiplier. 

The is income spending multiplier is  "=\frac{1}{1−MPC}"

Using the given information, 

income spending multiplier "=\frac{1}{1−MPC}"

income spending multiplier "=\frac{1}{1−0.75}"

income spending multiplier "=4"

"∆Y=Multiplier × ∆G \\ ∆Y=4 ×40 million\\∆Y= 160 million"

So, the change in output is 160 million.


(ii) The change in output from tax hike can be calculated with the help of tax multiplier.

Tax multiplier is "=\frac{-MPC}{1−MPC}"


Tax multiplier is"=\frac{-0.75}{1−0.75}"

Tax multiplier is"=\frac{-0.75}{0.25}"

Tax multiplier is"-3"

"∆Y= −3×∆T\\∆Y=−3×40 million\\∆Y= −120 million"

So, change in output from tax hike is -120 million.